OpenStudy (anonymous):

In the following equation, identify the x-intercepts in the graph y = 3x2 + 19x – 40.

hartnn (hartnn):

to find x-intercept put y=0,so u need to solve that quadratic equation...can u?

OpenStudy (anonymous):

0=6+19x-40?

hartnn (hartnn):

3x^2+19x-40=0....can u solve?

OpenStudy (anonymous):

3x^2+19x-40=0 -40 -40 3x^2+19x= -40

hartnn (hartnn):

u know quadratic equation formula?-b+root(b*b-4ac)/2a...this formula?

OpenStudy (anonymous):

nope lol

OpenStudy (campbell_st):

the quadratic will factorise....what are the factors of 40....

OpenStudy (campbell_st):

the factorised form is (3x + 8)(x - 5) = 0 you now need to find the intercepts by solving 3x + 8 = 0 and x - 5 = 0 good luck

hartnn (hartnn):

ae u sure campbell....coz i m getting (x+8)and (3x-5)

hartnn (hartnn):

which gives -8 and 5/3 and these are x-intercepts...